OpenStudy (anonymous):

Jim's uncle is 4 times as old as his nephew Jim. In 10 years, Jim's uncle's age will be 20 years more than twice Jim's age. How old is Jim's uncle right now?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

2(x+10)+20= 4x+10 2x+20+20=4x+10 -2x=-30 x=15

OpenStudy (anonymous):

jim = x uncle = 4x